OpenStudy (anonymous):

two numbers are in the ratio of 5:7.If 12 is added to each number the ratio is 4:5.What are the original numbers?

OpenStudy (anonymous):

Let the numbers be \(p\) , \(q\) . Then we have, \(\frac{p}{q} = \frac{5}{7}\) and \(\frac{p+12}{q+12} = \frac{4}{5} \) Solve the equations to get the actual numbers.